# Spectra and pseudospectra in the evaluation of material stability

### v2

@@ Abstract
We consider the dynamics of bodies with “active” microstructure described by vector-valued phase fields. For waves with time-varying amplitude, the associated evolution equation involves a matrix that can be non-normal, depending on the constitutive choices adopted for the microstructural actions associated with the phase field. In the presence of non-normality, the spectral analysis of such a matrix when (unknown or uncertain or time-dependent) constitutive parameters vary is generically not decisive to evaluate the material stability. We thus need to look at the pseudospectrum, that is, the set of all possible eigenvalues of matrices in an $\varepsilon$-neighborhood of the matrix of interest. As a case study, we look at quasicrystals and evaluate through spectra and pseudospectra the influence of a microstructural self-action on the material stability. This approach allows us to reliably determine thresholds where stable-unstable transitions in the material behavior occur.
@@ 1 Introduction
In the dynamics of bodies with “active” microstructure described by vector-valued phase fields we consider waves with time-varying amplitude. Their behavior is governed by the interactions between standard and microstructural actions, the latter ones associated with the (independent) mechanisms described by the phase field. Such microstructural interactions are a microstress (first-neighbor contact interaction) and a bulk self-action; their balance supplements the balance of macroscopic momentum and is coupled with it. No peculiar microstructural inertia [5] is attributed to phenomena described by the phase field; only possible microstructural (diffusion-type) dissipation is considered.
In this multi-field setting, constitutive structures may involve—and often they actually do—unknown or uncertain parameters. Parametric analyses may thus reveal admissible ranges for material stability and thresholds towards the unstable behavior. Results may thus address experiments aiming at the quantification of these unknown parameters. In developing such analyses in the mechanics of complex bodies (those with “active” microstructure) the occurrence of non-normal matrices is not infrequent, as in the case treated here of time-dependent amplitude waves. With the adjective “non-normal” we indicate those matrices $A\in\mathbb{C}^{n\times n}$ such that their product with the complex conjugate transpose matrix does not commute, namely $AA^{\ast}\neq A^{\ast}A$. For such matrices, the spectral analysis alone is in general not decisive for checking material stability or the transition to unstable behavior. Rather, we need to call upon the notion of pseudospectrum as introduced first by Trefethen and Trummer with reference to Chebyshev and Legendre spectral differentiation matrices, adopted in performing the explicit time (numerical) integration of an hyperbolic initial boundary value problem [37], [36] (see also [14]). The basic idea is to examine the spectra of matrices lying within a ball or radius $\varepsilon$ around the non-normal matrix $A$. The resulting sets of eigenvalues constitute the so-called pseudospectrum and provide more accurate information on possible stable-unstable transition in the mechanical behavior [36].
@@
As a concrete example, we perform explicit numerical analyses on wave propagation in quasicrystals. They are materials with perfect long-range order, but with no three-dimensional translation periodicity [32]. Their structure is made of atomic clusters distributed with a symmetry that is incompatible with a periodic tiling of the ambient space, for example icosahedral symmetry in $3D$-space [32], [31]. The quasi-periodic lattice is characterized by additional low-spatial-scale degrees of freedom, so-called phasons, here indicated by $\nu$ at $x$ and $t$. They are justified on geometrical basis. Indeed, if we expand in Fourier series the mass distribution of a $3D$ quasi-periodic lattice, the resulting wave vectors are six-dimensional [31]. More generally, a quasi-periodic lattice in a $n$-dimensional point space can be seen as a projection of a periodic lattice in a higher dimensional ambient onto a $nD$ incommensurate subspace. To visualize the circumstance, consider a $2D$ lattice with square symmetry; take in the plane a straight line inclined by an irrational angle with the lattice symmetry axes and project over it, orthogonally, the lattice atoms in a strip around the line (whatever be its thickness); the result is a quasi-periodic $1D$ lattice over the line, which is in this case the incommensurate subspace. A displacement in the higher-dimensional periodic lattice has thus a component in the incommensurate subspace and one orthogonal to it; the first projection of the displacement field is what is commonly called a phonon field, because it is associated with acoustic waves; the orthogonal displacement component to the incommensurate subspace leads to the field of phasons. This geometrical aspect suggested first a convenient description of the quasicrystal behavior under external actions as a higher-dimensional replica of the mechanics of simple bodies [12], [17]. However, if we consider the degrees of freedom represented by $\nu$ as those exploited by microscopic rearrangements that assure quasi-periodicity of the atomic lattice under given (possibly time-varying) boundary conditions, at continuum scale they are inner to every material element associated with a geometric point; so, $\nu$ itself has to be considered insensitive to relative rigid translations of observers in the physical space while, being a vector, it is sensitive to observer rotations. The circumstance detaches description of quasicrystal continuum mechanics from a higher dimensional replica of the standard format. Rather it falls within the general model-building framework for the mechanics of complex bodies (as shown in [22], [25]), in particular in the special case treated here, in which the phase field describing microstructural effects takes values in $\mathbb{R}^{3}$. In this general setting, balance equations are those of forces and those of microstructural interactions associated with the phase field; a balance of couples put together in the same expression standard and microstructural stresses, and a self-action at microstructure level. Although these balance equations can be postulated as first principles [4], they can be derived from an invariance principle together with the representation of contact actions in terms of standard and microstructural stresses. The principle mentioned can be the requirement of objectivity imposed to the external power alone [21] or the covariance (that is, structure-invariance under diffeomorphism-based observer changes) of the Clausius-Duhem inequality, written in terms of external power [24]. Resorting to first principles allows an appropriate control on the results unlike a mere use of analogies that can be surely useful but, in essence, are only hopes, although at times concrete ones.
@@
As regards quasicrystals, the invariance techniques mentioned (and here we adopt the $SO(3)$ invariance of the external power alone) imply the possible emergence of a phason self-action $\mathbf{z}$ which can have both conservative (energetic-type) and dissipative components [25]. In treatments of quasicrystal elasticity based on analogy to simple bodies in a higher dimensional spaces [12] a conservative component of $\mathbf{z}$ is completely discarded with the obvious advantage of reducing the analysis to a replica of standard elasticity with the proviso of considering a higher-dimensional space. When phason friction is considered, a dissipative component of $\mathbf{z}$ comes into play [30]. The conservative component of $\mathbf{z}$ should play, however, a role in material stability. To explore the circumstance, we analyze the time-dependent amplitude wave propagation in different cases:
@@
    purely dissipative self-action, so that $\mathbf{z}$ represents microstructural friction alone; thus, it is such that $\mathbf{z}\cdot\dot{\nu}\geq 0$, for any $\dot{\nu}$, an inequality that is compatible with the choice $\mathbf{z}=c\dot{\nu}$, $c\in\mathbb{R}^{+}$;
@@
    simultaneous presence of both conservative ($\mathbf{z}^{e}$) and dissipative ($\mathbf{z}^{d}$) components of the self-action, namely $\mathbf{z}=\mathbf{z}^{e}+\mathbf{z}^{d}$, where $\mathbf{z}^{e}=k_{0}\nu$ and $\mathbf{z}^{d}=c\dot{\nu}$.
@@
An interposed dot will eventually indicate the scalar product between elements of a linear space, whatever it be.
@@ 2 Basic fields
Consider two isomorphic copies of the $3D$ real space, namely ^$\hat{\mathbb{R}}^{3}$ and $\mathbb{R}^{3}$; the isomorphism $\hat{\iota}:\hat{\mathbb{R}}^{3}\longrightarrow\mathbb{R}^{3}$ is merely the identification map. In ^$\hat{\mathbb{R}}^{3}$ we select a fit region $\mathcal{B}$, a bounded simply connected domain endowed with a surface-like boundary oriented by the outward unit normal $\mathbf{n}$ to within a finite number of corners and edges. We adopt $\mathcal{B}$ as a macroscopic reference configuration. From it, deformations are orientation-preserving differentiable one-to-one maps $x\longmapsto y:=\tilde{y}(x)\in\mathbb{R}^{3}$. We indicate by $F$ the deformation gradient. To account for motions in a time interval $[0,\bar{t}\,]$, we extend $\tilde{y}$ to the space-time tube ×$\mathcal{B}\times[0,\bar{t}\,]$ so that we have $(x,t)\longmapsto y:=\tilde{y}(x,t)\in\mathbb{R}^{3}$ and we take $\tilde{y}(x,\cdot)$ to be differentiable also with respect to time at every $x\in\mathcal{B}$; a superposed dot will indicate the pertinent derivative. We thus define a displacement $\tilde{u}$ in the physical space as
@@
The velocity in Lagrangian representation is given by $\dot{y}=\frac{\mathrm{d}\tilde{y}(x,t)}{\mathrm{d}t}=\dot{u}$.
@@
brings at macroscopic scale information on microstructural events at spatial scale $\lambda$. ~$\tilde{\mathbb{R}}^{3}$ is isomorphic to $\mathbb{R}^{3}$ but distinguished from it; the isomorphism is once again the identification. We indicate by $N$ its gradient, namely $N:=\nabla\nu$. The time rate of $\nu$ in Lagrangian representation is given by $\dot{\nu}=\frac{\mathrm{d}\tilde{\nu}(x,t)}{\mathrm{d}t}$.
The small strain regime is commonly defined by the condition $|\nabla u|\ll 1$. In such a regime we do not distinguish between ^$\hat{\mathbb{R}}^{3}$ and $\mathbb{R}^{3}$.
@@ 3.1 Observers and their changes
The prescription of reference frames over all the spaces used for representing the morphology of a body and its motion is what we consider to be an observer. Here such spaces are the reference one (namely ^$\hat{\mathbb{R}}^{3}$), the physical space $\mathbb{R}^{3}$, ~$\tilde{\mathbb{R}}^{3}$ where $\tilde{\nu}$ takes values, and the time interval. We adopt observer changes that leave invariant the time scale and $\hat{\mathbb{R}^{3}}$, while they are isometries in $\mathbb{R}^{3}$ and ~$\tilde{\mathbb{R}}^{3}$.
More precisely, if $y$ is a place evaluated by a first observer $\mathcal{O}$, another observer $\mathcal{O}^{\prime}$ records a counterpart of $y$ given by $y^{\prime}=a(t)+Q(t)(y-y_{0})+y_{0}$, where $t\longmapsto a(t)\in\mathbb{R}^{3}$ is a vector-valued smooth map depending only on time; $y_{0}$ is an arbitrary fixed point, and $t\longmapsto Q(t)\in SO(3)$ is an orthogonal tensor-valued smooth map depending only on time.
@@
$\dot{y}^{\diamond}=\mathfrak{c}+q\times(y-y_{0})+\dot{y}\;.$
The vector $\mathfrak{c}:=Q^{\mathsf{T}}\dot{a}$ is a relative translation velocity between the two observers, while $q$ is the axial vector of the skew-symmetric second-rank tensor $Q^{\mathsf{T}}\dot{Q}$.
The vector $\nu:=\tilde{\nu}(x,t)\in\tilde{\mathbb{R}^{3}}$ collects degrees of freedom that are internal (thus relative) in time to the material element associated with $x\in\mathcal{B}$ in the reference configuration. This relative character renders $\nu$ a Galilean invariant vector, meaning it is insensitive to relative rigid translations between two generic observers. When $\mathcal{O}$ and $\mathcal{O}^{\prime}$ rotate relatively, the value $\nu$ recorded by a first observer changes into $\nu^{\prime}=Q(t)\nu$ and the time rate $\dot{\nu}$ becomes $\dot{\nu}^{\prime}=\dot{\nu}+\dot{Q}\nu$. By pulling back $\dot{\nu}^{\prime}$ into $\mathcal{O}$, we obtain a new vector $\dot{\nu}^{\diamond}:=Q^{\mathsf{T}}\dot{\nu}^{\prime}$ given by

### v3

@@ Abstract
We consider the dynamics of bodies with ‘active’ microstructure described by vector-valued phase fields. For waves with time-varying amplitude, the associated evolution equation involves a matrix that can be non-normal, depending on the constitutive choices adopted for the microstructural actions associated with the considered phase field. The occurrence of non-normality requires to look at the pseudospectrum of the considered matrix, namely the set of all possible eigenvalues of matrices in a $\varepsilon$-neighborhood of the matrix itself, because the eigenvalues of non-normal matrices can be very sensitive to small perturbations and therefore the spectral analysis alone would not be sufficient to distinguish with certainty between stable and unstable behavior. We develop the relevant analyses in the case of quasicrystals for which the values of some constitutive parameters are not known or are uncertain from an experimental point of view, a circumstance suggesting parametric analyses. We find circumstances in which the pseudospectra obtained by means of the so-called structured perturbations predict instability when, instead, the spectral analysis indicates stability.
@@ 1 Introduction
We discuss the propagation of waves with variable amplitude in bodies with ‘active’ microstructure described by vector-valued phase fields. The adjective ‘active’ refers to bodies with microstructure endowed with inner (but observable) degrees of freedom with pertinent interactions not directly associated with those kinematic mechanisms that are commonly associated with the standard stress (namely, the crowding and shearing of material elements merely considered as black boxes). In particular, we investigate the influence of a microstructural self-action on the stability of the material. Non-normal matrices emerge along the analysis. Their occurrence requires supplementing the conventional spectral approach to stability with an analysis of pseudospectra of the operators involved to verify stability and possible transitions to unstable behavior with certainty. Indeed, there may be conditions (expecially when constitutive parameters are uncertain or time-varying) in which pseudospectra indicate instability while, on the contrary, the standard spectral analysis predicts stability. In order to show concretely this phenomenon, in this paper we specifically refer to quasicrystals, but the approach has general validity.
@@
Quasicrystals are materials with perfect long-range order, but with no three-dimensional translation periodicity [44]. Their structure is made of atomic clusters distributed with a symmetry that is incompatible with a periodic tiling of the ambient space, for example icosahedral symmetry in $3D$-space [44], [43]. The quasi-periodic lattice is characterized by additional low-spatial-scale degrees of freedom, so-called phasons, here indicated by $\nu$ at $x$ and $t$. They are justified on geometrical basis. Indeed, if we expand in Fourier series the mass distribution of a $3D$ quasi-periodic lattice, the resulting wave vectors are six-dimensional [43]. More generally, a quasi-periodic lattice in a $n$-dimensional point space can be seen as a projection of a periodic lattice in a higher dimensional ambient onto a $nD$ incommensurate subspace. To visualize the circumstance, consider a $2D$ lattice with square symmetry; take in the plane a straight line inclined by an irrational angle with the lattice symmetry axes and project over it, orthogonally, the lattice atoms in a strip around the line (whatever be its thickness); the result is a quasi-periodic $1D$ lattice over the line, which is in this case the incommensurate subspace. Thus, a displacement in the higher-dimensional periodic lattice has a component in the incommensurate subspace and one orthogonal to it; the first projection of the displacement field is what we commonly call a phonon field, because it is associated with acoustic waves; the orthogonal displacement component to the incommensurate subspace leads to the field of phasons. This geometrical aspect first suggested a convenient description of the quasicrystal behavior under external actions as a higher-dimensional replica of the mechanics of simple bodies [15], [22]. However, if we consider the degrees of freedom represented by $\nu$ as those exploited by microscopic rearrangements that assure quasi-periodicity of the atomic lattice under given (possibly time-varying) boundary conditions, at continuum scale they are inner to every material element associated with a geometric point; so, $\nu$ itself has to be considered insensitive to relative rigid translations of observers in the physical space while, being a vector, it is sensitive to observer rotations. On the other side, even if we do not accept such an interpretation, since phasons belong to the space orthogonal to the incommensurate subspace (the physical space) on which we project the hyperlattice (that is the lattice in the higher-dimensional space), for a mere geometrical reason phasons are insensitive to observer translations in the physical space. All this implies that the mechanics of quasicrystals is more than a replica of the mechanics of simple bodies rewritten in a higher dimensional space. Rather it falls within the general model-building framework for the mechanics of complex bodies (as shown in [30], [34]). In this general setting, balance equations are those of forces and those of microstructural interactions associated with the phase field; a balance of couples links standard and microstructural stresses, and a self-action at microstructure level. Although these balance equations can be postulated as first principles [4], they can be derived from invariance principles together with the representation of contact actions in terms of standard and microstructural stresses. If we consider even the fully dissipative setting, the principles mentioned are the requirement of objectivity for the external power alone [29] or the covariance (that is, structure-invariance under diffeomorphism-based observer changes) of the Clausius-Duhem inequality, written in terms of external power [33]. The first path does not include the case in which the phase field is scalar or pseudo-scalar, the latter path covers even such cases. Resorting to first principles allows an appropriate control on the results unlike a mere use of analogies that can be surely useful but, in essence, are only hopes, although at times concrete ones.
@@
As regards quasicrystals, the invariance techniques mentioned (and here we adopt the $SO(3)$ invariance of the external power alone) imply the possible emergence of a phason self-action $\mathbf{z}$ which can have both conservative (energetic-type) and dissipative components [34], the latter indicated as phason friction [42]. The conservative component of $\mathbf{z}$ plays a role in the stability of the material. To explore the circumstance, we analyze the time-dependent amplitude wave propagation in different cases:
@@
    purely dissipative self-action, so that $\mathbf{z}$ represents microstructural friction alone; thus, it is such that $\mathbf{z}\cdot\dot{\nu}\geq 0$, for any $\dot{\nu}$, an inequality that is compatible with the choice $\mathbf{z}=\varsigma\dot{\nu}$, $\varsigma\in\mathbb{R}^{+}$;
@@
    simultaneous presence of both conservative ($\mathbf{z}^{e}$) and dissipative ($\mathbf{z}^{d}$) components of the self-action, namely $\mathbf{z}=\mathbf{z}^{e}+\mathbf{z}^{d}$, where $\mathbf{z}^{e}=k_{0}\nu$ and $\mathbf{z}^{d}=\varsigma\dot{\nu}$.
@@
The notion of pseudospectrum emerged from analyses in fluid dynamics and was later adopted in various other fields, from laser physics, control theory and others. Such a method deals with the spectral analysis of non-normal matrices. Precisely, in general, with $A\in\mathbb{C}^{n\times n}$ a $n\times n$ matrix over the field of complex numbers, and $A^{\ast}$ its conjugate transpose, we say that $A$ is non-normal when $AA^{\ast}\neq A^{\ast}A$. We can also evaluate the degree of non-normality [14]. However, besides formal definitions, let us come back to physics.
To evaluate the stability of fluid flows, we commonly (i) linearize the balance equations about the laminar state and (ii) look at the spectrum of the linearized operator to check whether there are unstable eigenvalues (those with real part in the positive half-plane). The method applies safely to various cases of simple fluids such as Bénard convection and Taylor-Couette flow. However, for the plane Couette flow the approach predicts linear stability for all Reynolds numbers, but experiments show transition to turbulence at Reynolds numbers as low as 350. Another failure in the theoretical stability predictions occurs, for example, in the analysis of the plane Poiseuille flow. Indeed, although the eigenvalues of the linearized operator lie in the stable half-plane, transient energy grows [51]. Non-orthogonality of eigenstates, which is a manifestation of non-normality of the governing operators, occurs in nuclear scattering, electronic conductance, and wave propagation in disordered media [10]. Also, convection-diffusion problems (damage evolution is an example) can be highly non-normal as many of the problems involving mixed derivatives [40].
In these cases, reliable stability information can be obtained by consideration of the $\varepsilon$-pseudospectrum (see, e.g., [50], [49], [18]). A complex number $z\in\mathbb{C}$ is in the $\varepsilon$-pseudospectrum of a matrix or, more generally, an operator $A$ if it is in the spectrum of some perturbed operator $A+B$ with $||B||<\varepsilon$. Pseudospectral material techniques for fluids have been largely developed for the Navier-Stokes equations (see, e.g., [16], [24], [38]) and a few other problems [49]).
The spectral analysis of the acoustic tensor is the key ingredient for evaluating the material stability in solids. Such a tensor is symmetric in the traditional setting of continuum mechanics; the matrix of its components has only real entries. This is essentially due to the symmetry of the Cauchy stress. The property obviously remains when we homogenize simple lattices, which lead to a traditional continuum scheme. In both cases, the spectral analysis generically suffices to evaluate the material stability.
When we consider bodies with active microstructure, as already mentioned, phase fields enter into the representation of the body morphology and are intended as observable entities. Interactions are associated with these fields and satisfy appropriate balance equations. Accounting for these interactions generically implies a lack of symmetry for the Cauchy stress [4], [29]. We also obtain this scheme when we homogenize complex lattices in which we have internal degrees of freedom (think for example of a lattice in which every material point is a small rigid body free to rotate about its mass center; the homogenized continuum is a Cosserat’s one), so that the homogenized continuum is not a classical (Cauchy-type) one but involves phase fields describing the additional degrees of freedom (rotations in the Cosserat case) [32]. However, in the linearized setting, even in the presence of microstructural actions the Cauchy stress reacquires always the symmetry [33]. This aspect notwithstanding, admissible constitutive relations for the microstructural interactions may induce non-normality, as we show in this paper.
Thus, when non-normality occurs, due to microstructural interactions, a pseudospectral analysis allows a refined (and reliable) evaluation of the material stability. To the best of our knowledge, we adopt for the first time the pseudospectral analysis in the mechanics of solids. Parametric analyses due to uncertainty and lack of knowledge of the constitutive parameters show circumstancxes in which the psedudospectra indicate possible instability while, instead, the spectral analysis show stability. We also highlight the important role played by (real) structured perturbations in addition to the more commonly used generic (complex) perturbations.
The paper is organized as follows: in Section 2 we indicate the basic fields associated with the body morphology. In Section 3 we discuss the notion of observer and obtain the balance equations, including those of the microstructural interactions. We avoid merely writing them down as postulates; we derive them from a requirement of objectivity of the external power to show clearly the origin and type of the lack of symmetry of the Cauchy stress. In Section 4 we discuss constitutive structures. In Section 5 we analyze the propagation of waves with variable amplitude in the linearized setting. To prepare and clarify the pseudospectral analysis, in Section 6 we recall basic notions and properties of the $\varepsilon$-pseudospectrum. Numerical results are discussed in Section 7. Some coefficients used in numerical simulations are known from experimental data. For those that are experimentally unknown, we perform parametric analyses in the admissible ranges that are theoretically established.
@@
An interposed dot will indicate the scalar product between elements of a linear space, whatever it be. As a further matter of notation, for a generic vector $k\in\mathbb{R}^{n}$, the symbol $|k|$ will indicate as usual the modulus of $k$. For a generic matrix $A$ with entries $a_{ij}$, $\|A\|$ will indicate the value of a generic norm; $\|\bullet\|_{2}$ will be written for the 2-norm, so that $\|A\|_{2}=s_{max}(A)$, where $s_{max}(A)$ is the maximum singular value of $A$; the symbol $\|\bullet\|_{F}$ will be used for the Frobenius norm defined by $\|A\|_{F}=\sqrt{\mathrm{tr}(A^{\ast}A)}$, where $\mathrm{tr}$ is the trace operator. The summation over repeated indices is occasionally adopted throughout the paper.
@@ 2 Basic fields
Consider two isomorphic copies of the $3D$ real space, namely ^$\hat{\mathbb{R}}^{3}$ and $\mathbb{R}^{3}$; the isomorphism $\hat{\iota}:\hat{\mathbb{R}}^{3}\longrightarrow\mathbb{R}^{3}$ is merely the identification map. In ^$\hat{\mathbb{R}}^{3}$ we select a fit region $\mathfrak{B}$, a bounded simply connected domain endowed with a surface-like boundary oriented by the outward unit normal $\mathbf{n}$ to within a finite number of corners and edges. We adopt $\mathfrak{B}$ as a macroscopic reference configuration. From it, deformations are orientation-preserving differentiable one-to-one maps $x\longmapsto y:=\tilde{y}(x)\in\mathbb{R}^{3}$. We indicate by $F$ the deformation gradient. To account for motions in a time interval $[0,\bar{t}\,]$, we extend $\tilde{y}$ to the space-time tube $\mathfrak{B}\times[0,\bar{t}\,]$ so that we have $(x,t)\longmapsto y:=\tilde{y}(x,t)\in\mathbb{R}^{3}$ and we take $\tilde{y}(x,\cdot)$ to be differentiable also with respect to time at every $x\in\mathfrak{B}$; a superposed dot will indicate the time derivative. We thus define a displacement $\tilde{u}$ in the physical space as
@@
In Lagrangian representation the velocity is $\dot{y}=\frac{\partial\tilde{y}(x,t)}{\partial t}=\dot{u}$.
@@
brings at macroscopic scale information on microstructural events at spatial scale $\lambda$; its choice is matter of modeling, so it depends on the cases considered. In the special case of quasicrystals considered here, the ratio between the modulus of $\nu$ and the one of $u$ is about $10^{-2}$, as shown by numerical simulations in [35] on the basis of experimental data.
~$\tilde{\mathbb{R}}^{3}$ is isomorphic to $\mathbb{R}^{3}$ but distinguished from it; the isomorphism is once again the identification. We indicate by $N$ its spatial derivative, namely
$N:=\nabla\nu=\bigg{(}\frac{\partial\nu^{i}}{\partial x}\bigg{)}^{A}\tilde{\mathbf{e}}_{i}\otimes\mathbf{e}_{A}\,,$
where $\tilde{\mathbf{e}}_{i}$ is the $ith$ element of a dual basis over ~$\tilde{\mathbb{R}}^{3}$ and $\mathbf{e}^{A}$ is the $Ath$ element of the dual basis over ^$\hat{\mathbb{R}}^{3}$. The time rate of $\nu$ in Lagrangian representation is given by $\dot{\nu}=\frac{\mathrm{d}\tilde{\nu}(x,t)}{\mathrm{d}t}$.
The small strain regime is commonly defined by the condition $\|\nabla u\|_{F}\ll 1$. In such a regime we do not distinguish between ^$\hat{\mathbb{R}}^{3}$ and $\mathbb{R}^{3}$.
@@ 3.1 Observers and their changes
The prescription of reference frames over all the spaces used for representing the morphology of a body and its motion is what we consider to be an observer. Here such spaces are the reference one (namely ^$\hat{\mathbb{R}}^{3}$), the physical space $\mathbb{R}^{3}$, the one where $\tilde{\nu}$ takes values, namely ~$\tilde{\mathbb{R}}^{3}$, and the time interval. We adopt observer changes that leave invariant the time scale and $\hat{\mathbb{R}^{3}}$, while they are isometries in $\mathbb{R}^{3}$, with consequent changes of reference frames on ~$\tilde{\mathbb{R}}^{3}$.
More precisely, if $y$ is a place evaluated by a first observer $\mathcal{O}$, another observer $\mathcal{O}^{\prime}$ records a counterpart of $y$ given by $y^{\prime}=a(t)+Q(t)(y-y_{0})+y_{0}$, where $t\longmapsto a(t)\in\mathbb{R}^{3}$ is a time-dependent vector-valued smooth map; $y_{0}$ is an arbitrary fixed point, and $t\longmapsto Q(t)\in SO(3)$ is a time-dependent orthogonal tensor-valued smooth map.
@@
$\dot{y}^{\diamond}=c+q\times(y-y_{0})+\dot{y}\;.$
The vector $c:=Q^{\mathsf{T}}\dot{a}$ is a relative translation velocity between the two observers, while $q$ is the axial vector of the skew-symmetric second-rank tensor $Q^{\mathsf{T}}\dot{Q}$.
The vector $\nu:=\tilde{\nu}(x,t)\in\tilde{\mathbb{R}^{3}}$ collects degrees of freedom that are internal (thus relative) in time to the material element associated with $x\in\mathfrak{B}$ in the reference configuration. Thus $\nu$ is a Galilean invariant vector, meaning it is insensitive to relative rigid translations between two generic observers in the physical space. When $\mathcal{O}$ and $\mathcal{O}^{\prime}$ rotate relatively, the value $\nu$ recorded by a first observer changes into $\nu^{\prime}=Q(t)\nu$ and the time rate $\dot{\nu}$ becomes $\dot{\nu}^{\prime}=\dot{\nu}+\dot{Q}\nu$. By pulling back $\dot{\nu}^{\prime}$ into $\mathcal{O}$, we obtain a new vector $\dot{\nu}^{\diamond}:=Q^{\mathsf{T}}\dot{\nu}^{\prime}$ given by